The momentum of a body increases by 20%. What is the % increase it`s K.E ?

Grade:12

3 Answers

Ganesh Domale
6 years ago
KE increases by 400%KE=(p^2)/2m You have given ratios of momentum..Take ratios of initial and final KE emergy and get your answer
Hamza
6 years ago
(mv2-mv1)÷mv1=20/100mv2/mv1=(1/5+1)v1/v2=5/6(Ke1/Ke2)=(v1/v2)^2=25/36(Ke2/ke1-1)=(36/25-1)=11/25Percent=(11/25)×100=44%
Hamza
6 years ago
Plz verify my answer it is correct u cant disapprove it(mv2-mv1)÷mv1=20/100mv2/mv1=(1/5+1)v1/v2=5/6(Ke1/Ke2)=(v1/v2)^2=25/36(Ke2/ke1-1)=(36/25-1)=11/25Percent=(11/25)×100=44%